OpenStudy (anonymous):

The expression 4x/5 - 2x/3 is equivalent to... a) x/2 b) x/5 c) x/15 d) 2x/15

OpenStudy (greatlife44):

now remember to subtract fractions we need to multiply both by the LCD

OpenStudy (greatlife44):

\[\frac{ 4x }{ 5 }*(15)-\frac{ 2x }{ 3 }*(15)\]

OpenStudy (greatlife44):

Which results in the following \[\frac{ (4x*3) -(2x*5)}{ 15 }\]

OpenStudy (anonymous):

I'm confused. What would be the final answer then? Sorry, I don't really get subtracting fractions.

OpenStudy (greatlife44):

well, we need to make the denominator the same if you notice they are different 3 and 5, so we must multiply each fraction by the least common denominator of 3 and 5 which is 15. this gives us \[\frac{ 2x }{ 15 }\]

OpenStudy (anonymous):

Thank you so much! Would you mind helping me with other problems if you have time?